OpenStudy (anonymous):

write the fraction as percent 19/25

OpenStudy (jamesj):

Ok tankie, this is your 3rd or 4th percentage question You should begin to figure this out by yourself.

OpenStudy (jamesj):

What is 45/100 as a percent? or 63/100? In short, you should know by now how to handle things like something/100

OpenStudy (jamesj):

So the question then is: how do you get this expression 9/25 to be something/100

OpenStudy (jamesj):

If 9/25 = something/100, what is something?

OpenStudy (jamesj):

You've gone quiet ... so I'll help \[\frac{9}{25} = \frac{4 \times 9}{4 \times 25} = \frac{36}{100}\]

OpenStudy (anonymous):

\[\frac{19}{25} = \frac{4 \times 19}{4 \times 25} = \frac{76}{100} = 76%\]